### v2

@@ 5.1 Data and Methodology
I use the China Annual Survey of Industrial Firms (CASIF), a longitudinal micro-level dataset collected by the National Bureau of Statistics of China that includes information on all state-owned industrial firms and non-state-owned firms with annual sales above 5 million RMB (~US\$770K). According to Brandt, Van Biesebroeck, Wang, and Zhang (2017), they account for 91% of the gross output, 71% of employment, 97% of exports, and 91% of total fixed assets in 2004, and thus are representative of industrial activities in China. Many papers on topics such as firm behavior, international trade, and growth theory have used the CASIF data (e.g., Hsieh and Klenow (2009), Brandt, Van Biesebroeck, Wang, and Zhang (2017)).
I focus on the five largest 2-digit sectors in terms of the number of firms between 2004 and 2007.222222There are 3140 textile firms, 2186 chemical firms, 1804 nonmetallic minerals firms, 2869 general equipment firms, and 1568 transportation equipment firms in the final dataset. I order the industries by their 2-digit Chinese Industry Classification codes. [^23] The CASIF dataset spans between 1998 and 2007. I choose year 2004 to 2007 to (i) ensure data consistency due to the change in the Chinese Industry Classification codes in 2003, (ii) avoid the financial crisis in the early 2000s, and (iii) use the most recent data. Following Brandt, Van Biesebroeck, and Zhang (2014), appropriate price deflators for inputs and outputs are applied separately. I preprocess the data so that firms with strictly positive amounts of capital, employment, value-added output, real wage expense, and real interest rate are used for estimation. The final dataset consists of a balanced panel of 11,567 firms over four years across five sectors. Summary statistics of the key variables are presented in Table 1.
@@
[…]
@@
The value-added production function I consider is
$y_{it}$ $=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
$\beta_{it}^{K}$ $=\beta^{K}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \beta_{it}^{L}=\beta^{L}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \omega_{it}=\omega\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),$
$k_{it}$ $=g^{K}\left(Z_{it},A_{i},\eta_{it}^{K}\right),\ l_{it}=g^{L}\left(Z_{it},A_{i},\eta_{it}^{L}\right).$ (5.1)
@@
I highlight two features of model (5.1). First, output elasticities $\beta_{it}\coloneqq\left(\beta_{it}^{K},\beta_{it}^{L}\right)^{{}^{\prime}}$ are allowed to differ across firms and through time. Second, input choices $X_{it}\coloneqq\left(k_{it},l_{it}\right)^{{}^{\prime}}$ can be correlated with $\beta_{it}$ through their dependence on the pairwise correlated random variables $A_{i}$, $\eta_{it}$, and $\varepsilon_{it}$.
It is worth noting that the output is measured by the total revenue in dollars, rather than by the physical quantity produced in units, due to the absence of individual output prices in the data. When firms operate in distinct imperfectly competitive output markets, this may cause problems (Klette and Griliches (1996)). I use input prices $Z_{it}\coloneqq\left(r_{it},w_{it}\right)^{{}^{\prime}}$ as IVs for $X_{it}$. [^24] Real wages are very likely to be exogenous because the labor market in the manufacturing sectors of China is close to a perfectly competitive market in which firms face a relatively flat supply curve. Real interest rate may be considered as a valid IV because its fluctuation is mostly driven by monetary policy set by the central bank of China. For estimation, I take $W_{i}$ to be the mean through time of each coordinate of $X_{it}$. I estimate each coordinate of $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$ by regressing $\mathbf{\mathbbm{1}}\left(k\leq k_{it}\right)$ and $\mathbf{\mathbbm{1}}\left(l\leq l_{it}\right)$ on the second-degree polynomial spline basis of $\left(Z_{it},W_{i}\right)$ with its knot at the median, respectively. Next, I estimate $G_{it}\coloneqq\mathbb{E}\left[\left.y_{it}\right|X_{it},V_{it},W_{i}\right]$ by regressing $y_{it}$ on the same spline basis of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$. Estimation of $b_{1t}\left(V_{it},W_{i}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ is then obtained by taking the partial derivative of $\widehat{G}_{it}\left(X_{it},\widehat{V}_{it}W_{i}\right)$ with respect to $X_{it}$. Finally, I estimate $\overline{b}\coloneqq\mathbb{E}\beta_{it}$ by averaging $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ over $i$ and $t$ and $b_{t}\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by regressing $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the second-degree polynomial spline basis of $X_{it}$ with its knot at the median.
@@ 5.2 Results
First, I summarize the results for estimating the average output elasticities defined as $\widehat{\overline{b}}\coloneqq\left(\widehat{\overline{b}}^{K},\widehat{\overline{b}}^{L}\right)^{{}^{\prime}}$ in Table 2. I compare my estimated $\widehat{\overline{b}}$’s (TERC) with those derived by applying the methods of Olley and Pakes (1996) (OP), Levinsohn and Petrin (2003) (LP), and Ackerberg, Caves, and Frazer (2015) (ACF) to the same data set. Standard errors (se) and 95% confidence intervals (CI) are also included. [^25] I use the Stata commands prodest (Rovigatti and Mollisi (2018)) for OP and LP and acfest (Manjón and Manez (2016)) for ACF for point estimates of $\overline{b}$ and 95% CI. See Keiller, de Paula, and Van Reenen (2024) for a discussion of the implementation of the ACF method. For TERC, since $\widehat{\overline{b}}$ has slower than $n^{1/2}$ rate by Theorem 2, I follow Politis, Romano, and Wolf (1999) to randomly subsample $b=\lfloor 4n^{3/4}\rfloor$ firms from each sector without replacement for 1,000 times.
@@
Capital Elasticity 0.359 0.252 0.300 0.387
se (0.030) (0.023) (0.037) (0.031)
95% CI [0.300, 0.417] [0.206, 0.297] [0.227, 0.373] [0.334, 0.455]
Labor Elasticity 0.470 0.175 0.567 0.375
se (0.014) (0.011) (0.065) (0.023)
95% CI [0.442, 0.498] [0.154, 0.196] [0.440, 0.695] [0.327, 0.418]
@@
Capital Elasticity 0.294 0.288 0.344 0.380
se (0.036) (0.031) (0.096) (0.037)
95% CI [0.223, 0.365] [0.228, 0.348] [0.155, 0.533] [0.306, 0.451]
Labor Elasticity 0.296 0.113 0.378 0.329
se (0.022) (0.015) (0.213) (0.026)
95% CI [0.253, 0.339] [0.084, 0.143] [-0.039, 0.795] [0.276, 0.377]
@@
Capital Elasticity 0.697 0.311 0.236 0.379
se (0.101) (0.024) (0.101) (0.037)
95% CI [0.499, 0.895] [0.263, 0.358] [0.038, 0.434] [0.314, 0.451]
Labor Elasticity 0.353 0.071 0.601 0.459
se (0.021) (0.010) (0.384) (0.027)
95% CI [0.311, 0.394] [0.051, 0.091] [-0.151, 1.354] [0.404, 0.514]
@@
Capital Elasticity 0.416 0.246 0.176 0.504
se (0.076) (0.023) (0.082) (0.033)
95% CI [0.267, 0.565] [0.202, 0.291] [0.016, 0.337] [0.451, 0.578]
Labor Elasticity 0.444 0.071 0.927 0.359
se (0.019) (0.009) (0.149) (0.020)
95% CI [0.406, 0.482] [0.053, 0.089] [0.635, 1.218] [0.320, 0.397]
@@
Capital Elasticity 0.523 0.281 0.217 0.383
se (0.064) (0.033) (0.084) (0.035)
95% CI [0.397, 0.649] [0.216, 0.346] [0.052, 0.381] [0.327, 0.469]
Labor Elasticity 0.523 0.137 1.042 0.460
se (0.025) (0.019) (0.138) (0.028)
95% CI [0.473, 0.573] [0.100, 0.173] [0.771, 1.313] [0.407, 0.517]
My estimated average capital elasticities are within the range of $\left[0.379,0.504\right]$ across the five sectors, which are consistent with the results obtained from applying OP, LP, and ACF to the same dataset. The estimated average labor elasticities present more discrepancies between methods. My estimates of $\widehat{\overline{b}}^{L}$’s lie between $\left[0.329,0.460\right]$ across the five sectors, similar to OP’s results. LP seem to derive small $\widehat{\overline{b}}^{L}$’s for all sectors, while ACF generate larger $\widehat{\overline{b}}^{L}$’s for the general equipment sector and transportation equipment sector. My 95% CI’s for $\widehat{\overline{b}}$’s are reasonably tight across all sectors. It is well documented in the literature that output elasticities in Cobb-Douglas production function estimation usually lie within $\left[0,1\right]$. Hsieh and Klenow (2009) also calculate that roughly half of output is distributed to capital, according to Chinese input-output tables and national accounts. My estimates are consistent with the empirical evidence in the literature.
@@
Next, I investigate the distribution of the conditional means of the output elasticities given the regressors. Policymakers could use the information contained in the conditional means to formulate more accurate sector specific policies, as these conditional means provide a fine approximation to the average elasticities for each subgroup of firms with a certain level of capital and labor. Moreover, distributional properties about the true output elasticities can be inferred from the distribution of the conditional means of the output elasticities. For example, by the law of total variance, the variance of the conditional means of the output elasticities provides a lower bound for the variance of the true output elasticities. Demirer (2022) argues that the heterogeneity in output elasticities is largely explained by across-firm variation. To investigate into this explanation, for each firm I calculate the averages of $\widehat{b}_{t}^{K}\left(X_{it}\right)$ and $\widehat{b}_{t}^{L}\left(X_{it}\right)$ through time and denote them by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$, respectively. Then, I plot the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across firms for each of the five sectors.
@@
In Figure 2, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the textile sector and indicate the corresponding $\widehat{\overline{b}}$’s labeled by “Mean” on the same graph. The left subplot of Figure 2 shows the histogram of $\widehat{\beta}_{i}^{K}$. All of its probability mass lies between zero and one, with its mode at 0.4. Over 90% of the probability mass of $\widehat{\beta}_{i}^{K}$ lies between 0.3 and 0.5. The right subplot of Figure 2 presents the histogram of $\widehat{\beta}_{i}^{L}$ across firms for the textile sector. Again, the majority of its probability mass lies between zero and one, with its mode at 0.38. The distributions of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ are concentrated around the mean, suggesting that firms in the textile industry tend to be homogeneous in terms of capital and labor efficiency.
@@
In Figure 3, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the other four sectors. First, the majority of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across all sectors lie between zero and one, which is consistent with empirical evidence. Second, I find significant across-firm variations in both capital and labor elasticities measured by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ within each of the five sectors. However, the degree of heterogeneity in these elasticities differs. For example, in the left subplot of Figure 3, there is a larger variance in $\widehat{\beta}_{i}^{K}$ in the nonmetallic mineral sector than in the chemical sector, which implies that the firms in the chemical industry may be more homogeneous in capital efficiency than those in the nonmetallic mineral sector.

### v3

@@ 5.1 Data and Methodology
I use the China Annual Survey of Industrial Firms (CASIF), a longitudinal micro-level dataset collected by the National Bureau of Statistics of China that includes information on all state-owned industrial firms and non-state-owned firms with annual sales above 5 million RMB (~US\$770K). According to Brandt, Van Biesebroeck, Wang, and Zhang (2017), they account for 91% of the gross output, 71% of employment, 97% of exports, and 91% of total fixed assets in 2004, and thus are representative of industrial activities in China. Many papers on topics such as firm behavior, international trade, and growth theory have used the CASIF data (e.g., Hsieh and Klenow (2009)).
I focus on the five largest 2-digit sectors in terms of the number of firms between 2004 and 2007. Note that the CASIF dataset spans between 1998 and 2007. I choose year 2004 to 2007 to (i) ensure data consistency due to the change in the Chinese Industry Classification codes in 2003, (ii) avoid major structural breaks in the early 2000s (e.g., China joined the WTO in 2001), and (iii) use the most recent data.
Following Brandt, Van Biesebroeck, and Zhang (2014), appropriate price deflators for inputs and outputs are applied separately. I preprocess the data so that firms with strictly positive amounts of capital, employment, value-added output, real wage expense, and real interest rate are used for estimation. The final dataset consists of a balanced panel of 11,317 firms over four years across five sectors. Summary statistics are presented in Table 1.
@@
$y_{it}=\ln$$\left(\text{value-added output}\right)$ 45,268 9.55 1.34 2.24 16.96
$k_{it}=\ln$$\left(\text{capital}\right)$ 45,268 9.17 1.56 0.98 16.84
$l_{it}=\ln$$\left(\text{labor}\right)$ 45,268 5.08 1.08 2.08 11.97
$Z_{it,1}=\ln\overline{r}_{cs(i),t}^{\text{real},(-i)}$ 45,268 0.73 0.45 -6.02 6.94
$Z_{it,2}=\ln\overline{w}_{cs(i),t}^{\text{real},(-i)}$ 45,268 2.68 0.37 0.33 5.28
@@
  • Notes:
    (i) Output is measured as firm-level real value added in constant 1998 RMB (10,000 yuan). Capital is measured as firm-level real fixed capital stock (book value net of depreciation), deflated to constant 1998 RMB (10,000 yuan). Labor is measured as the number of employees (persons). The real interest rate is defined as real interest expenditures divided by total debt (percent). The real wage is average real salary per worker per year in constant 1998 RMB (10,000 yuan).
    (ii) Sample sizes by industry are: textile (3,072), chemicals (2,131), nonmetallic mineral products (1,758), general equipment (2,828), and transportation equipment (1,528).
    (iii) $Z_{it,1}$ denotes the natural logarithm of debt-weighted average real interest rate of competitors in the same city–industry–year. $Z_{it,2}$ denotes the natural logarithm of employment-weighted average real wage of competitors in the same city–industry–year.
The value-added heterogeneous production function is
$y_{it}$ $=k_{it}\beta_{it,K}+l_{it}\beta_{it,L}+\omega_{it}+\widetilde{\epsilon}_{it},$
$\beta_{it,K}$ $=\beta_{K}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \beta_{it,L}=\beta_{L}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \omega_{it}=\omega\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),$
$k_{it}$ $=g_{K}\left(Z_{it},A_{i},\eta_{it,K}\right),\ l_{it}=g_{L}\left(Z_{it},A_{i},\eta_{it,L}\right).$ (5.1)
@@
I highlight two features of model (5.1). First, output elasticities $\beta_{it}\coloneqq(\beta_{it,K},\beta_{it,L})^{\prime}$ are allowed to differ across firms and through time. Second, input choices $X_{it}\coloneqq(k_{it},l_{it})^{\prime}$ can be correlated with $\beta_{it}$ through their dependence on the pairwise correlated random variables $A_{i}$, $\eta_{it}$, and $\varepsilon_{it}$.
I use leave-one-out industry–city–year cell averages of real wages and real interest rates faced by other firms in the same industry and city, weighted by competitors’ employment (for wages) and total debt (for interest rates) as IVs. I write them as $Z_{it}\coloneqq(\ln\overline{r}_{cs(i),t}^{\text{real},(-i)},\,\ln\overline{w}_{cs(i),t}^{\text{real},(-i)})^{\prime}$, where $cs(i)$ represents city and sector of firm $i$. These instruments are relevant because firms compete locally for labor and capital, so variation in competitors’ cost conditions shifts firm $i$’s effective input prices and hence its optimal input choices. At the same time, they are plausibly exogenous to firm $i$’s idiosyncratic productivity shocks and fixed effects because firm $i$’s own input prices are excluded from the construction; identification therefore relies on competitors’ cost variation rather than firm-level wage or interest-rate realizations. In Appendix B, I examine the exogeneity issue by using a more exogenous IV constructed at the provincial level, and the results remain stable.
For estimation, I take $W_{i}$ to be the mean through time of each coordinate of $X_{it}$ by adapting the argument of Mundlak (1978) nonparametrically. I also examine the effect of further including the mean through time of each coordinate of $Z_{it}$ and the results reported in Table A.3 are similar. For all series estimation steps, I use second-degree polynomial spline basis functions with knots at the median, and a robustness check to this choice is included in Appendix B. First, I estimate each coordinate of $V_{it}\coloneqq(F_{\left.k_{it}\right|Z_{it},W_{i}}(\left.k_{it}\right|Z_{it},W_{i}),F_{\left.l_{it}\right|Z_{it},W_{i}}(\left.l_{it}\right|Z_{it},W_{i}))^{\prime}$ by regressing $\mathbf{\mathbbm{1}}(k_{it}\leq k)$ and $\mathbf{\mathbbm{1}}(l_{it}\leq l)$ on the basis functions of $(Z_{it},W_{i})$, respectively. Next, I estimate $G_{it}\coloneqq\mathbb{E}[\left.y_{it}\right|X_{it},V_{it},W_{i}]$ by regressing $y_{it}$ on the basis functions of $(X_{it},\widehat{V}_{it},W_{i})$. Estimation of $b_{1t}(V_{it},W_{i})\coloneqq\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]$ is then obtained by taking the partial derivative of $\widehat{G}_{it}(X_{it},\widehat{V}_{it},W_{i})$ with respect to $X_{it}$. Finally, I estimate the pooled APE $\overline{b}$ by simply averaging $\widehat{b}_{1t}(\widehat{V}_{it},W_{i})$ over $i$ and $t$. For the LAR, as in (4.7), I regress $\widehat{b}_{1t}(\widehat{V}_{it},W_{i})$ on the basis functions of $X_{it}$ for each $t$, then average over $t$ to obtain $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ as a measure of heterogeneity in output elasticities across firms.
@@ 5.2 Results
Average Output Elasticities
I compare my TERC estimates of average output elasticities with those obtained from OP (Olley and Pakes (1996)), LP (Levinsohn and Petrin (2003)), and ACF (Ackerberg, Caves, and Frazer (2015)) applied to the same dataset. Table 2 contains the estimation results.
@@
Capital Elasticity 0.359 0.252 0.300 0.415
95% CI [0.300, 0.417] [0.206, 0.297] [0.227, 0.373] [0.393, 0.441]
Labor Elasticity 0.470 0.175 0.567 0.452
95% CI [0.442, 0.498] [0.154, 0.196] [0.440, 0.695] [0.415, 0.485]
@@
Capital Elasticity 0.294 0.288 0.344 0.463
95% CI [0.223, 0.365] [0.228, 0.348] [0.155, 0.533] [0.430, 0.497]
Labor Elasticity 0.296 0.113 0.378 0.311
95% CI [0.253, 0.339] [0.084, 0.143] [-0.039, 0.795] [0.263, 0.361]
@@
Capital Elasticity 0.697 0.311 0.236 0.371
95% CI [0.499, 0.895] [0.263, 0.358] [0.038, 0.434] [0.332, 0.410]
Labor Elasticity 0.353 0.071 0.601 0.311
95% CI [0.311, 0.394] [0.051, 0.091] [-0.151, 1.354] [0.255, 0.365]
@@
Capital Elasticity 0.416 0.246 0.176 0.433
95% CI [0.267, 0.565] [0.202, 0.291] [0.016, 0.337] [0.402, 0.469]
Labor Elasticity 0.444 0.071 0.927 0.521
95% CI [0.406, 0.482] [0.053, 0.089] [0.635, 1.218] [0.479, 0.564]
@@
Capital Elasticity 0.523 0.281 0.217 0.425
95% CI [0.397, 0.649] [0.216, 0.346] [0.052, 0.381] [0.393, 0.459]
Labor Elasticity 0.523 0.137 1.042 0.588
95% CI [0.473, 0.573] [0.100, 0.173] [0.771, 1.313] [0.536, 0.644]
  • Notes: For TERC, I implement the three-step series estimator as described above. To compute the CI, I use a subsampling method to draw $b=\lfloor 4n_{s}^{3/4}\rfloor$ firms from each sector without replacement and repeat this 1,000 times. For the other methods, I use the Stata commands (prodest, Rovigatti and Mollisi (2018)) for OP and LP and (acfest, Manjón and Manez (2016)) for ACF.
The $\widehat{\overline{b}}_{K}$ estimates from TERC fall within $\left[0.371,\,0.463\right]$ across the five sectors, broadly consistent with the results obtained from applying OP, LP, and ACF to the same dataset. In contrast, the $\widehat{\overline{b}}_{L}$ estimates exhibit greater discrepancies across methods. The TERC estimates of $\widehat{\overline{b}}_{L}$ lie in $\left[0.311,0.588\right]$ across the five sectors. OP’s labor elasticity estimates are close to TERC’s for the first three sectors but are smaller for the last two. LP seems to produce smaller $\widehat{\overline{b}}_{L}$ estimates across all sectors, whereas ACF generates larger $\widehat{\overline{b}}_{L}$ estimates for the general equipment and transportation equipment sectors. It is worth noting that the ACF approach tends to produce wider confidence intervals in certain sectors, which may partly reflect numerical instability in the Stata implementation, an issue also discussed by Keiller, de Paula, and Van Reenen (2024). The 95% CIs for TERC estimates are reasonably tight across all sectors.
I consider the estimation results from TERC reasonable based on two pieces of empirical evidence. First, it is well documented in the literature that output elasticities in Cobb-Douglas production function estimation usually lie within $\left[0,1\right]$. Second, using Chinese manufacturing data, Hsieh and Klenow (2009) show that roughly half of output accrues to capital. Under profit maximization with a Cobb-Douglas production function, output elasticities correspond to factor revenue shares. The magnitudes of the estimated capital and labor elasticities can therefore be interpreted through this lens, and my estimates are consistent with this result.
The broad similarity of average elasticities across methods is not unexpected, as OP, LP, ACF, and the TERC approach are all designed to address simultaneity between input choices and unobserved productivity. While OP, LP, and ACF provide consistent estimates under a constant-coefficient production function, TERC adds value by relaxing this restriction and allowing output elasticities to vary across firms while still addressing simultaneity via a control-function approach.
Remark 4 (Addressing Simultaneity Issue).
OP, LP, and ACF address simultaneity via a proxy-variable strategy: they invert a strictly monotone proxy input demand—investment in OP and intermediate inputs in LP/ACF—to recover a scalar productivity term $\omega_{it}$, then impose a Markov-style law of motion for $\omega_{it}$ to form identifying moments. This requires (i) strict monotonicity in $\omega_{it}$ and (ii) that $\omega_{it}$ is the only econometric unobservable entering the proxy demand, which rules out random coefficients that directly affect firms’ input choices. In contrast, my baseline TERC model allows input choices to correlate with heterogeneous output elasticities through an arbitrary-dimensional fixed effect $A_{i}$ and a time-varying shock $\eta_{it}$, with productivity modeled flexibly as $\omega_{it}=\omega(A_{i},\varepsilon_{it})$. Identification proceeds via a control-function strategy using a sufficient statistic $W_{i}$ for $A_{i}$ and a period-specific control $V_{it}$ for $\eta_{it}$, so that conditional on $(V_{it},W_{i})$ the remaining input variation is driven by exogenous instruments without taking a Markov law for $\omega_{it}$ as a primitive assumption.
Cross-Firm Heterogeneity in Output Elasticities
@@
Next, I examine the distribution of conditional mean output elasticities given the regressors. These conditional means summarize average elasticities for firm subgroups defined by their capital and labor levels, and thus provide a useful basis for designing more targeted, sector-specific policies. They are also informative about the underlying distribution of true elasticities. For instance, by the law of total variance, the variance of these conditional means is a lower bound on the variance of the true output elasticities. Demirer (2025) argues that the heterogeneity in output elasticities is largely explained by across-firm variation. To study this issue, I average $\widehat{b}_{t,K}(X_{it})$ and $\widehat{b}_{t,L}(X_{it})$ for each firm through time and denote them by $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$, respectively. These $\widehat{\beta}_{i}$ estimates can be considered as a proxy for their output elasticities. Then, I plot the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ across firms for each sector.
In Figure 1, I present the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ for the textile sector and indicate the corresponding pooled APE labeled by “Mean” in the figure. The left subplot of Figure 1 corresponds to the histogram of $\widehat{\beta}_{i,K}$. All of its probability mass lies between zero and one, with its mode at 0.4. Over 95% of its probability mass lies between 0.3 and 0.5. The distribution of $\widehat{\beta}_{i,K}$ is concentrated around the mean and symmetric, suggesting that these firms tend to be homogeneous in capital efficiency. The right subplot of Figure 1 is the histogram of $\widehat{\beta}_{i,L}$. Again, the majority of its probability mass lies between zero and one. The distribution of $\widehat{\beta}_{i,L}$ is more dispersed than that of $\widehat{\beta}_{i,K}$ and is slightly left-skewed, with a small number of textile firms exhibiting low labor efficiency.
@@
Figure 2 presents the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ for the other sectors. I draw the following two conclusions. First, for all sectors, the majority of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ lie between zero and one, consistent with the empirical evidence in Hsieh and Klenow (2009). Second, there is substantial across-firm variations in both capital and labor elasticities within each sector. The extent of heterogeneity, however, differs across elasticities and sectors. For example, in the left panel of Figure 2, capital elasticity exhibits greater dispersion among chemical firms than among firms in the nonmetallic minerals sector.
@@
Explaining the Heterogeneity
I examine cross-firm heterogeneity in output elasticities by regressing the estimated $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ on firm characteristics. Specifically, using the CASIF dataset, I construct five covariates: firm size, leverage, export status, industrial-park location, and state ownership. All specifications include industry and city fixed effects, and I report heteroskedasticity-robust standard errors.
Table 3: Explaining the Heterogeneity
$\widehat{\beta}^{K}$ $\widehat{\beta}^{L}$
Firm Size 0.0328 (0.0006) 0.0002 (0.0012)
Leverage 0.0004 (0.0024) 0.0261 (0.0044)
Export Status 0.0098 (0.0010) 0.0023 (0.0018)
Industrial Park -0.0014 (0.0013) -0.0043 (0.0026)
State Ownership 0.0050 (0.0040) -0.0534 (0.0076)
  • Notes:
    (i) Size is defined as the natural logarithm of total average assets. Leverage is measured as the debt-to-asset ratio. Exporting firms are defined as those with positive export delivery value. Industrial park firms are identified based on firm address information containing keywords such as “development zone” or “industrial park.” State-owned enterprises (SOEs) are defined as firms whose state capital accounts for more than 50 percent of total paid-in capital.
    (ii) The variables are measured as follows: levels for the elasticities; the natural logarithm of 10,000 yuan for firm size; a ratio for leverage; and binary indicators (0/1) for the last three variables in the table.
    (iii) The numbers in parentheses in the second row of each cell represent the standard errors.
The results summarized in Table 3 show that larger firms have significantly higher capital elasticities, while leverage is positively associated with labor elasticities. Exporters exhibit modestly higher capital elasticities, and state-owned enterprises have substantially lower labor elasticities. Conditional on these controls, the industrial-park indicator is small. Industry and location fixed effects are jointly significant, indicating that sectoral and regional factors account for an important share of the heterogeneity. The adjusted-$R^{2}$s are 0.58 and 0.57 for $\widehat{\beta}^{K}$ and $\widehat{\beta}^{L}$, respectively, suggesting that over half of the variation in estimated elasticities can be explained by observable firm characteristics and fixed effects.
The results are generally intuitive and broadly consistent with existing empirical evidence. Larger firms, for instance, often operate in more capital-intensive segments or adopt technologies with stronger scale economies, automation, and standardized production processes, so output tends to be more responsive to capital at the margin (Syverson (2011)). Leverage may also be associated with more labor-intensive operating choices—for example, delaying or outsourcing capital investment, relying on older or rented equipment, or placing greater emphasis on variable inputs (Kalemli-Özcan, Laeven, and Moreno (2022)). To the extent that highly leveraged firms substitute away from capital toward labor, the estimated production relationship will place greater weight on labor. Finally, exporters tend to use capital more efficiently (Bernard and Jensen (1999)), whereas state-owned enterprises may employ labor in excess of the efficient level due to political considerations (Wen (2025)), which is consistent with their substantially lower labor elasticities relative to private firms.